My son has this problem on his math. I don't have any idea where to start. We really just need help figuring out what the equation is. We can take it from there.
"There is a game at a fair in which you toss balls into a bowl to win fish. There are 52 bowls. The larger bowls contain 5 fish, and the smaller ones contain 3 fish. If there are 202 fish total, how many bowls of each size are there?"

Found 3 solutions by scott8148, Earlsdon, pavel111:
Answer by scott8148(6628)

it is a system of equations ___ two unknowns, two equations

"There are 52 bowls" ___ L + S = 52

"The larger bowls contain 5 fish, and the smaller ones contain 3 fish. If there are 202 fish total"
___ 5L + 3S = 202

solve by substitution or elimination

Answer by Earlsdon(6294)

First, let L = the number of large bowls and S = the number of small bowls. From the problem statement, you have:
1) L+S = 52 "There are 52 bowls."
The number of fish in a large bowl = 5 while a smaller bowl has 3 fish. This is represented algebraically as:
5*L and 3*S and the sum of all the fish is 202, so...
2) 5L+3S = 202
Rewrite equation 1) as:
1a) L = 52-S and substitute into equation 2).
2a) 5(52-S)+3S = 202 Now you can solve for S.
2b) 260-5S+3S = 202
2c) 260-2S = 202 Subtract 260 from both sides.
2d) -2S = -58 Finally, divide both sides by -2.
2e) S = 29 and...
L = 52-S
L = 52-29
L = 23
There are 23 large bowls and 29 small bowls.

Answer by pavel111(1)